OpenStudy (anonymous):

why is 48 the answer for f(x) varies directly with x2, and f(x) = 75 when x = 5, find the value of f(4)

OpenStudy (campbell_st):

so with direct variation the equation is \[f(x) = k \times x^2\] so use the given information to find k... \[75 = k \times 5^2\] when you have k the equation becomes \[f(4) = k \times 4^2\]